OpenStudy (anonymous):

Derivative of : (1-t^4)/(2) Explain it please step by step ...

OpenStudy (anonymous):

f(x)'=2t^3

OpenStudy (anonymous):

step by step :D ?

OpenStudy (anonymous):

f(x)=a/b , f(x)'=(a'b-b'a)/b^2

OpenStudy (anonymous):

OHHHH ;D thanks

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

\[\frac{1-t^4}{2}=\frac{1}{2}-\frac{t^4}{2}\] \[\frac{d}{dt}[\frac{1}{2}-\frac{t^4}{2}]=-{2t^3}\]

OpenStudy (anonymous):

correct